# The neutrino force at all length scales

### v1

@@ Abstract
The Standard Model predicts a long-range force mediated by a pair of neutrinos, known as “the neutrino force.” We derive an expression for this force that is valid for all $r$. For large $r$, it reduces to the known $G_{F}^{2}/r^{5}$ form, while for small $r$, it scales as $1/r$. We explore the implications of this result for atomic parity violation (APV) experiments. A key feature of the neutrino force is that it is a long-range effect, meaning it cannot be treated as a correction to the tree-level exchange $Z$ diagram. We calculate the effects in muonium and positronium, finding that the neutrino force contributes about 4% and 16%, respectively, compared to the leading $Z$ exchange. This indicates a significant impact on APV, with important implications for detecting the neutrino force and measuring the weak mixing angle in APV experiments.
@@ 1 Introduction
Beyond macroscopic probes, the search for the neutrino force extends into atomic and subatomic scales Dzuba:2017cas ; Ghosh:2019dmi . One promising phenomenon that probes this force is atomic parity violation (APV) (see e.g. Safronova:2017xyt ; Wieman:2019vik ; Arcadi:2019uif for a review). The neutrino force violates parity and contributes to the forbidden electric-dipole transition element in the atom. As a result, it can be probed by looking at the parity-violating atomic transition matrix element. Ref. Ghosh:2019dmi calculated the neutrino-induced APV effects for the hydrogen atom in the four-Fermi effective theory. However, the calculation in Ghosh:2019dmi is only valid for $\ell\geq 2$ (where $\ell$ is the orbital angular momentum), the result of which is too small to be probed. For smaller $\ell$, the calculations of the atomic matrix element do not converge, indicating the failure of four-Fermi approximation. In APV experiments, so far, measurements were done for $s$-wave states (i.e., $\ell=0$). The wave function of such states is concentrated at $r=0$, a region where the four-Fermi approximation breaks down. Therefore, in order to perform a self-consistent calculation of the APV effects caused by neutrinos for $\ell=0$, it is necessary to know the short-range behavior of the neutrino force. This calls for a calculation of APV effects using a form for neutrino force that is valid at all length scales.
@@
The rest of the paper is organized as followed. In Sec. 2, we briefly review the neutrino force in the framework of the four-Fermi effective theory and explain the limitations of the result. In Sec. 3, we present the calculation of the neutrino force in the full electroweak theory, including all the relevant diagrams. The implications of our results to experiments are discussed in the next three sections. In Sec. 4, we introduce the basic idea to use APV as an experimental probe of the neutrino force. In Sec. 5, we calculate the effects of neutrino forces on APV observables in muonium and positronium. We also highlight the important impact of the neutrino force on extracting the weak mixing angle from APV experiments. Other quantum forces in the SM as well as their APV effects are calculated in Sec. 6 for completeness. We comment on and compare our results with existing literature in Sec. 7 and conclude in Sec. 8. Some technical details of our computations are provided in the two appendices.
@@ 2 The neutrino force in the four-Fermi effective theory
$\mathcal{L}_{\rm eff}=\frac{G_{F}}{\sqrt{2}}\left[\overline{\chi}\gamma^{\mu}(g_{V}^{\chi}-g_{A}^{\chi}\gamma^{5})\chi\right]\left[\overline{\nu}\gamma_{\mu}(1-\gamma^{5})\nu\right],$ (2)
@@ 3.1 Formalism
where $\mathcal{A}_{\text{NR}}\equiv\mathcal{M}/(4m_{1}m_{2})$ denotes the normalized amplitude in the NR limit and $q=p_{1}-p_{1}^{\prime}$ is the momentum transfer. In the NR limit we have $q\approx(0,{\bf q})$.
The integral in Eq. (7) can be computed using the dispersion technique. See Ref. Feinberg:1989ps for a detailed review of the dispersion technique and its application to calculating the long-range forces. For instance, the spin-independent potential is given by
@@
is the discontinuity of the spin-independent amplitude in the complex $t$-plane. To get ${\cal A}_{\rm NR}(t)$, one should replace $q^{2}$ with $t$ in the amplitude. Although the amplitude ${\cal A}_{\rm NR}$ calculated from quantum field theories may be divergent, the discontinuity part of the one-loop amplitude is always finite according to the optical theorem. Therefore, the result in Eq. (8) is also always finite.
In the most general case, there are three different components of the neutrino forces in Eq. (7): the spin-independent (SI) part, the spin-dependent parity-conserving (SD-PC) part, and the spin-dependent parity-violating (PV) part. Note that the spin-independent part is always parity-conserving and the parity-violating part must be spin-dependent. In this paper, we focus on the SI part and the PV part of the force, where the latter contributes to the APV experiments. We work in the Pauli-Dirac basis in the NR limit, where the wave-functions have the following solution:
@@ 3.2.1 The self-energy diagram
In the long-distance limit, $r\gg 1/m_{Z}$, we have
@@ 3.2.2 The penguin diagrams
There are two penguin diagrams of concern, one with a vertex correction at the fermion $\chi_{1}$, as shown in Fig. 2(b), and another diagram with a vertex correction at the anti-fermion $\overline{\chi_{2}}$ as in Fig. 2(c). The potential $V_{0}$ for two penguin diagrams has same structure except for the dependence on the external fermion mass $m_{i}$. At leading order of $m_{i}^{2}/m_{Z}^{2}$ we have:
$V_{0}^{\rm PG}(r)=-\frac{g^{4}}{2048\pi^{3}c_{W}^{2}r}\int_{0}^{\infty}{\rm d}t\,e^{-\sqrt{t}\,r}\,\frac{t\left(2m_{W}^{2}+3t\right)-2\left(m_{W}^{2}+t\right)^{2}\log\left(1+t/{m_{W}^{2}}\right)}{\left(t-4m_{i}^{2}\right)^{2}\left(t-m_{Z}^{2}\right)}\;,$ (21)
where $i=1,2$ denotes the contribution from Fig. 2(b) and 2(c), respectively. The integral in Eq. (21) cannot be computed analytically but can be calculated numerically at any distance, as shown in Fig. 3(b). Note that the fermion mass $m_{i}$ in the denominator is crucial to cancel the infrared divergence of the integral.
In the long-distance limit, $r\gg m_{Z}^{-1}$, Eq. (21) reduces to the familiar $1/r^{5}$ form:
@@
In the short-distance limit $r\ll m_{Z}^{-1}$, we have
$V_{0}^{\rm PG}(r)$ $=\frac{g^{4}}{2048\pi^{3}c_{W}^{2}}\frac{1}{r}\left\{4\log\left(m_{W}r\right)\left[2\gamma_{\rm E}+\log\left(m_{W}r\right)\right]+6\log\left(m_{i}r\right)\right.$
$\left.\qquad\qquad\qquad\quad+\pi^{2}+2\gamma_{\rm E}\left(3+2\gamma_{\rm E}\right)+3\left(1+2\log 2\right)\right\}$
@@
where in the last step we have used the fact that at sufficiently short distances, $\log^{2}(m_{W}r)$ dominates over $\log(m_{W}r)$ and $\log(m_{i}r)$. Therefore, the dependence of $V_{0}^{\rm PG}$ on the external fermion mass drops out in both the long-distance and the short-distance limits.
@@ 3.2.3 The box diagram
$V_{0}^{\rm box}(r)$ $=-\frac{g^{4}}{1024\pi^{3}r}\left[4\log^{2}\left(m_{W}r\right)+8\gamma_{\rm E}\log\left(m_{W}r\right)+4\gamma_{\rm E}^{2}+\pi^{2}\right]$
@@ 3.2.4 Combination of all diagrams
To have a better understanding of the evolution of the neutrino force with distances, we showed $V_{0}$ for each diagram in Figs. 3 and 4, based on Eqs. (17), (21) and (24). (For $V_{0}^{\rm PG}$ in the figures, we have fixed $m_{i}$ to be the electron mass for definiteness.)
@@ 4.1 Review of APV
However, previous studies have neglected the effects of neutrinos on APV. For instance, Refs. Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss calculated the one-loop corrections to APV in the SM up to ${\cal O}(\alpha G_{F})$. The neutrino one-loop effects, which are ${\cal O}(G_{F}^{2})$, were ignored as they were assumed to be too small. This is correct at the amplitude level, because all amplitudes involving neutrino loops in Fig. 2 vanish in the limit of $q^{2}\to 0$, i.e., for the neutrino one-loop contribution we get $\kappa_{1}=\kappa_{2}=0$ in Eq. (42). This limit is equivalent to taking $m_{Z}^{2}\to\infty$ at the amplitude level, namely, taking the $Z$ force as a $\delta$-function. However, as illustrated in Fig. 9, the neutrino force has a much longer range than the $Z$ force. As a result, taking the contact-interaction limit at the amplitude level cannot capture the long-range effects of neutrinos on APV. Therefore, the weak charge defined in Eq. (42) is not an appropriate quantity to describe APV when there exists long-range parity-violating potentials.
To improve upon the existing results, one needs to take into account the full distribution of the parity-violating neutrino force in the atomic system and convolve it with the atomic wavefunctions. In fact, as we will show below, after including the full-range expressions, the generic neutrino force contribution to the atomic parity-violating matrix elements for $s$-wave states ($\ell=0$) is of order $g^{2}/(16\pi^{2})\sim 0.3\%$ when compared with the leading $Z$ effect. Therefore, the neutrino contribution to APV is at the same order as the ${\cal O}(\alpha G_{F})$ contributions computed in Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss , and should be included for a complete calculation of the one-loop radiative corrections to the parity-violating matrix elements in atoms. Remarkably, the contribution from pure neutrinos is comparable to both the current APV experimental error ($\sim$ 0.35%) Wood:1997zq and the theoretical uncertainty in the atomic many-body systems calculations ($\sim$ 0.45%) Porsev:2009pr ; Dzuba:2012kx . This makes us optimistic about the use of atomic systems in probing the neutrino force in the not too distant future.
@@
  • ($ii$)
    For states with angular momentum $\ell>0$, the electron wavefunctions are not concentrated at the origin, making the tree-level contribution highly suppressed by the very short range of the $Z$ force. In this case, the neutrino force contribution can even dominate over the $Z$ force.
Note that these two enhancements effects are not realized in current APV experiments, which have only been done in heavy atoms for $s$-wave states.
@@ 5.1 Parity-violating forces in muonium and positronium systems
where $V_{0}^{\rm SE}$ and $V_{0}^{\rm PG}$ are given by Eqs. (17) and (21), respectively. The factors in front of each $V_{0}$ are determined by the effective couplings $G_{V}G_{A}$ in Tab. 1, see Eq. (14). In particular, there are two penguins diagrams giving the factor $(+1)(-1)+(+1)(4s_{W}^{2}-1)=-(2-4s_{W}^{2})$. Note here we have neglected the difference in $V_{0}^{\rm PG}$ caused by different external fermion masses (namely, $m_{i}=m_{e}$ or $m_{i}=m_{\mu}$) so that the spin dependence in Eq. (50) can be factored out. The external fermion mass $m_{i}$ is much smaller than the mass of the gauge bosons, so it can be completely neglected at long distances. At short distances, the dependence of $V_{0}^{\rm PG}$ on $m_{i}$ is logarithmic and is much smaller compared to the squared logarithm of $m_{W}$, as shown in Eq. (3.2.2), so it does not cause sizable difference between the case of $m_{i}=m_{e}$ and the case of $m_{i}=m_{\mu}$. Numerically, we also confirm that at all ranges $V_{0}^{\rm PG}$ is not sensitive to external fermion masses as long as they are much smaller than the gauge boson mass.
@@ 5.2.1 $\ell=0$ state
So far, the APV measurements have only been done for $s$-wave states ($\ell=0$). For concreteness, we consider the transition matrix element between $s$-wave and $p$-wave states, i.e, schematically $|A\rangle=|1S\rangle$ and $|B\rangle=|2P\rangle$. (We have checked that the numerical value of the ratio $\eta_{\nu}$ does not change significantly for other small values of the principal quantum number.) Note that the energy difference in the denominator of Eq. (45) shall cancel in this ratio and we can focus on the matrix elements $\langle 2P|V^{j}_{\rm PV}|1S\rangle$ alone.
@@
An important observation regarding the $s$-wave states that simplifies our efforts greatly is that the $s$-wave wavefunctions do not change greatly with respect to the parity violating potential. This, in turn allows the wave functions and their derivatives to be treated as roughly constant in the region where the potential $V_{0}$ is defined. In this approximation the spin operator contributions [see Eqs. (47)-(48) and (50)-(51)] to $\eta$ can be cancelled in the numerator and denominator leaving only the radial integrals to be performed. Note that this simplification can only be performed in the case of $s$-waves. Neglecting the spin effects for states with higher angular momentum is not advisable. Nevertheless, we provide as an estimate for the case of $p$-wave and $d$-wave states later in the paper [see Eqs. (66)-(67)]. For a complete treatment of the spin operators for $\ell\geq 2$ states with four-Fermi approximation, see Ghosh:2019dmi .
@@
5.2.2 $-\ell$ states
For $-\ell$ states, the atomic wavefunctions are not concentrated at the origin, thus the tree-level $Z$ contribution is heavily suppressed, thereby increasing the ratio $\eta_{\nu}$. For instance, using the same strategy above, for the self-energy diagram, we obtain:

### v2

# Neutrino force at all length scales

@@ Abstract
The Standard Model predicts a long-range force mediated by a pair of neutrinos, known as “the neutrino force”. It scales as $G_{F}^{2}/r^{5}$, where $G_{F}$ is the Fermi constant. However, as $r\lesssim\sqrt{G_{F}}$, the four-Fermi theory breaks down and the neutrino force no longer has the $1/r^{5}$ scaling. For the first time, we derive a complete expression for the neutrino force that is valid at all distances. For $r\gg\sqrt{G_{F}}$, the result reduces to the known $G_{F}^{2}/r^{5}$; for $r\ll\sqrt{G_{F}}$, it scales as $1/r$. We explore the implications of this result for atomic parity violation (APV) experiments. A key feature of the neutrino force is that it is a long-range effect compared to the atomic length scale. Thus, in general, it cannot be simply treated as a correction to the tree-level $Z$-exchange diagram without considering the atomic wavefunctions. We calculate the effects in muonium and positronium, finding that the neutrino force contributes about 4% and 16%, respectively, compared to the leading $Z$ exchange. This indicates a significant impact on APV, with important implications for detecting the neutrino force and measuring the weak mixing angle in APV experiments.
@@ 1 Introduction
Beyond macroscopic probes, the search for the neutrino force extends into atomic and subatomic scales Dzuba:2017cas ; Ghosh:2019dmi . One promising phenomenon that probes this force is atomic parity violation (APV) (see e.g. Safronova:2017xyt ; Wieman:2019vik ; Arcadi:2019uif for a review). The neutrino force violates parity and contributes to the forbidden electric-dipole transition element in the atom. As a result, it can be probed by looking at the parity-violating atomic transition matrix element. Ref. Ghosh:2019dmi calculated the neutrino-induced APV effects for the hydrogen atom in the four-Fermi effective theory. However, the calculation in Ghosh:2019dmi is only valid for $\ell\geq 2$ (where $\ell$ is the orbital angular momentum), the result of which is too small to be probed. For smaller $\ell$, the calculations of the atomic matrix element do not converge, indicating the failure of four-Fermi approximation. In APV experiments, so far, measurements were done for $s$-states (i.e., $\ell=0$). The wave function of such states is concentrated at $r=0$, a region where the four-Fermi approximation breaks down. Therefore, in order to perform a self-consistent calculation of the APV effects caused by neutrinos for $\ell=0$, it is necessary to know the short-range behavior of the neutrino force. This calls for a calculation of APV effects using a form for neutrino force that is valid at all length scales.
@@
The rest of the paper is organized as followed. In Sec. 2, we briefly review the neutrino force in the framework of the four-Fermi effective theory and explain the limitations of the result. In Sec. 3, we present the calculation of the neutrino force in the full electroweak theory, including all the relevant diagrams. The implications of our results to experiments are discussed in the next three sections. In Sec. 4, we introduce the basic idea to use APV as an experimental probe of the neutrino force. In Sec. 5, we calculate the effects of neutrino forces on APV observables in muonium and positronium. We also highlight the important impact of the neutrino force on extracting the weak mixing angle from APV experiments. Other quantum forces in the SM as well as their APV effects are calculated in Sec. 6 for completeness. We comment on and compare our results with existing literature in Sec. 7 and conclude in Sec. 8. Some technical details of our computations are provided in the three appendices.
@@ 2 The neutrino force in the four-Fermi effective theory
$\mathcal{L}_{\rm eff}=-\frac{G_{F}}{\sqrt{2}}\left[\overline{\chi}\gamma^{\mu}(g_{V}^{\chi}-g_{A}^{\chi}\gamma^{5})\chi\right]\left[\overline{\nu}\gamma_{\mu}(1-\gamma^{5})\nu\right],$ (2)
@@ 3.1 Formalism
where $\mathcal{A}_{\text{NR}}\equiv\mathcal{M}/(4m_{1}m_{2})$ denotes the normalized amplitude in the NR limit and $q=p_{1}-p_{1}^{\prime}$ is the momentum transfer. In the NR limit, we have $q\approx(0,{\bf q})$.
The integral in Eq. (7) can be computed using the dispersion technique. See Ref. Feinberg:1989ps for a detailed review of the dispersion technique and its application to calculate long-range forces. (The original formalism developed in Ref. Feinberg:1989ps assumed that there was no additional pole on the real $t$-axis apart from the branch cut. In Appendix C, we generalize the dispersion formalism in the presence of additional poles, which is relevant for the calculation of the neutrino force from self-energy and penguin diagrams.)
Using the dispersion formalism, the spin-independent potential is given by
@@
is the discontinuity of the spin-independent amplitude in the complex $t$-plane. Here, $t$ is the value of $q^{2}$ after analytic continuation (see Appendix C for details). Although the amplitude ${\cal A}_{\rm NR}$ calculated from quantum field theories may be divergent, the discontinuity part of the one-loop amplitude is always finite according to the optical theorem. Therefore, the result in Eq. (8) is also always finite. As we show below, the parity-violating potentials can be also be obtained from Eq. (8) with some spin-dependent prefactors.
In the most general case, there are three different components of the neutrino forces in Eq. (7): the spin-independent (SI) part, the spin-dependent parity-conserving (SD-PC) part, and the spin-dependent parity-violating (PV) part Ghosh:2024qai . Note that the spin-independent part is always parity-conserving and the parity-violating part must be spin-dependent. In this paper, we focus on the SI part and the PV part of the force, where the latter contributes to the APV experiments. We work in the Pauli-Dirac basis in the NR limit, where the wave-functions have the following solution:
@@ 3.2.1 The self-energy diagram
In Fig. 3(a), $V_{0}^{\rm SE}$ is plotted as a function of the distance.
In the long-distance limit, $r\gg 1/m_{Z}$, the expression in Eq. (17) is reduced to
@@ 3.2.2 The penguin diagrams
There are two penguin diagrams of concern, one with a vertex correction at the fermion $\chi_{1}$, as shown in Fig. 2(b), and another diagram with a vertex correction at the anti-fermion $\overline{\chi_{2}}$ as in Fig. 2(c). The potential $V_{0}$ for two penguin diagrams has the same structure except for the dependence on the external fermion mass $m_{i}$. At leading order of $m_{i}^{2}/m_{Z}^{2}$ we have:
$V_{0}^{\rm PG}(r)=-\frac{g^{4}}{2048\pi^{3}c_{W}^{2}r}\int_{0}^{\infty}{\rm d}t\,e^{-\sqrt{t}\,r}\,\frac{t\left(2m_{W}^{2}+3t\right)-2\left(m_{W}^{2}+t\right)^{2}\log\left(1+t/{m_{W}^{2}}\right)}{t^{2}\left(t-m_{Z}^{2}\right)}\;,$ (21)
where the dependence on the fermion mass drops out.
The $1/(t-m_{Z}^{2})$ factor in Eq. (21) comes from the $Z$ propagator in the penguin diagram, which brings about a singularity for the integral. To obtain a finite and physical result, the pole should be shifted by an infinitesimally small imaginary part according to the Feynman propagator prescription: $t-m_{Z}^{2}\to t-m_{Z}^{2}+i\epsilon$. In the limit of $\epsilon\to 0^{+}$, a finite imaginary part appears due to the identity ${\rm Im}\left[1/(t-m_{Z}^{2}+i0^{+})\right]=-\pi\delta(t-m_{Z}^{2})$. However, as we show in Appendix C, this imaginary part is unphysical and is exactly canceled by the imaginary part of the residue of the amplitude at $Z$ pole. As a result, the final expression for the interaction potential is real, as it should be. Therefore, practically, the integral in Eq. (21) should be computed in the following way:
$V_{0}^{\rm PG}(r)$ $=-\frac{g^{4}}{2048\pi^{3}c_{W}^{2}r}$
$\times{\rm Re}\left[\lim_{\epsilon\to 0^{+}}\int_{0}^{\infty}{\rm d}t\,e^{-\sqrt{t}\,r}\,\frac{t\left(2m_{W}^{2}+3t\right)-2\left(m_{W}^{2}+t\right)^{2}\log\left(1+t/{m_{W}^{2}}\right)}{t^{2}\left(t-m_{Z}^{2}+i\epsilon\right)}\right].$ (22)
The integral in Eq. (3.2.2) is well defined and can be calculated numerically at any distance. The result is shown in Fig. 3(b).
The analytical expressions for $V_{0}^{\rm PG}$ can be found in the asymptotic regions. In the long-distance limit, $r\gg m_{Z}^{-1}$, Eq. (21) reduces to the familiar $1/r^{5}$ form:
@@
In the short-distance limit $r\ll m_{Z}^{-1}$, using a Mellin transform, we get
$V_{0}^{\rm PG}(r)$ $=\frac{g^{4}}{2048\pi^{3}c_{W}^{2}r}\left[4\log^{2}\left(e^{\gamma_{\rm E}}m_{W}r\right)+6\log\left(e^{\gamma_{\rm E}}m_{Z}r\right)-\log^{2}\left(\frac{m_{W}^{2}}{m_{Z}^{2}}\right)+\frac{4\pi^{2}}{3}\right],$
@@
where in the last step we have used the fact that at sufficiently short distances, $\log^{2}(m_{W}r)$ dominates over all other terms.
@@ 3.2.3 The box diagram
Note that this integral is well defined and can be numerically computed at arbitrary distances, as shown in Fig. 3(c).
@@
$V_{0}^{\rm box}(r)$ $=-\frac{g^{4}}{1024\pi^{3}r}\left[4\log^{2}\left(e^{\gamma_{\rm E}}m_{W}r\right)+\frac{\pi^{2}}{2}-1\right]$
@@ 3.2.4 Combination of all diagrams
To have a better understanding of the evolution of the neutrino force with distances, we showed $V_{0}$ for each diagram in Figs. 3 and 4, based on Eqs. (17), (3.2.2) and (25).
@@ 4.1 Review of APV
However, previous studies have neglected the effects of neutrinos on APV. For instance, Refs. Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss calculated the one-loop corrections to APV in the SM up to ${\cal O}(\alpha G_{F})$. The effects of two-neutrino exchange, which are of the order ${\cal O}(G_{F}^{2})$, were ignored because they were assumed to be too small. The calculations in Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss were performed only at the amplitude level in the limit of $q^{2}\to 0$ without taking into account the radial dependence of the atomic wavefunctions. This method can only capture the APV effect when the atomic wavefunctions are concentrated around the origin and is not generically applicable when there exist long-range parity-violating potentials.
To improve upon the existing results, one needs to take into account the full distribution of the parity-violating neutrino force in the atomic system and convolve it with the atomic wavefunctions. This is precisely described by the accumulated charge defined in Eq. (71) and plotted in Fig. 5. For $s$-states ($\ell=0$), since the atomic wavefunctions are roughly constant around the origin, this convolution is essentially the integral of the neutrino force itself over the atomic length scale; for $\ell\geq 1$ states, the radial dependence of the atomic wavefunctions plays a nontrivial role in the convolution. Phenomenologically, the current most promising observable for APV is the $s$-state parity-violating matrix elements in heavy atoms Wood:1997zq . As we will show below, after correctly taking into account the short-range behavior of the neutrino force, the generic contribution from two-neutrino exchange to the atomic parity-violating matrix elements for $s$-states is of order $g^{2}/(16\pi^{2})\sim 0.3\%$ when compared with the leading $Z$ effect. Therefore, the neutrino contribution to APV is at the same order as the ${\cal O}(\alpha G_{F})$ contributions computed in Marciano:1978ed ; Marciano:1982mm ; Marciano:1983ss , and should be included for a complete calculation of the one-loop radiative corrections to the parity-violating matrix elements in atoms. Remarkably, the contribution from pure neutrinos is comparable to both the current APV experimental error ($\sim$ 0.35%) Wood:1997zq and the theoretical uncertainty in the atomic many-body systems calculations ($\sim$ 0.45%) Porsev:2009pr ; Dzuba:2012kx . This makes us optimistic about the use of atomic systems in probing the neutrino force in the not too distant future.
@@
  • (iiii)
    For states with angular momentum $\ell>0$, in the non-relativistic approximation, the electron wavefunctions are not concentrated at the origin, making the tree-level contribution highly suppressed by the very short range of the $Z$ force. In this case, the relative contribution of the neutrino force is enhanced.
    (In heavy atoms, the relativistic corrections from the Dirac equation, as well as the effects from electron correlations, are non-negligible, which increases the electron density around the origin for $\ell>0$ states. As a result, the relative enhancement of the neutrino contribution is suppressed.)
Note that these two enhancement effects are not realized in current APV experiments, which have been done only in heavy atoms for $s$-states.
@@ 5.1 Parity-violating forces in muonium and positronium systems
where $V_{0}^{\rm SE}$ and $V_{0}^{\rm PG}$ are given by Eqs. (17) and (21), respectively. The factors in front of each $V_{0}$ are determined by the effective couplings $G_{V}G_{A}$ in Tab. 1, see Eq. (14). In particular, there are two penguins diagrams giving the factor $(+1)(-1)+(+1)(4s_{W}^{2}-1)=-(2-4s_{W}^{2})$.
@@ 5.2.1 $\ell=0$ state
So far, the APV measurements have only been done for $s$-states ($\ell=0$). For concreteness, we consider the transition matrix element between the $s$- and $p$-states, i.e, schematically $|A\rangle=|1S\rangle$ and $|B\rangle=|2P\rangle$. (We have checked that the numerical value of the ratio $\eta_{\nu}$ does not change significantly for other small values of the principal quantum number.) Note that the energy difference in the denominator of Eq. (46) shall cancel in this ratio and we can focus on the matrix elements $\langle 2P|V^{j}_{\rm PV}|1S\rangle$ alone.
@@
An important observation regarding the $s$-states that simplifies our efforts greatly is that the $s$-state wavefunctions do not change greatly with respect to the parity-violating potential. This, in turn allows the wave functions and their derivatives to be treated as roughly constant in the region where the potential $V_{0}$ is defined. In this approximation, the spin operator contributions [see Eqs. (48)-(49) and (51)-(52)] to $\eta$ can be canceled in the numerator and denominator, leaving only the radial integrals to be performed. Note that this simplification can only be performed in the case of $s$-states. Neglecting the spin effects for states with higher angular momentum is not advisable. Nevertheless, we provide as an estimate for the case of $p$-states and $d$-states later in the paper [see Eqs. (67)-(68)]. For a complete treatment of the spin operators for $\ell\geq 2$ states with four-Fermi approximation, see Ghosh:2019dmi .
@@
To obtain Eqs. (59)-(60) from Eqs. (56)-(57), we have changed the order of integrations — we first perform the radial part integral $\int{\rm d}r$ and then perform the momentum square integral $\int{\rm d}t$. Using this trick, although the potentials $V_{0}^{\rm PG}$ and $V_{0}^{\rm box}$ in Eqs. (21) and (25) do not have analytical expressions, the atomic matrix elements induced by them can be computed analytically, as shown in Eqs. (59)-(60). We have also verified that they agree very well with the direct result obtained from 2D numerical integrals.
@@
5.2.2 Higher-$\ell$ states
For higher-$\ell$ states, in the non-relativistic limit, the atomic wavefunctions are not concentrated at the origin, thus the tree-level $Z$ contribution is heavily suppressed, thereby increasing the ratio $\eta_{\nu}$. For instance, using the same strategy above, for the self-energy diagram, we obtain: